Nitrogen gas having a mass of 0.8 kg at 110 kPa and 50℃ is contained in a cylinder. The piston is moved compressing nitrogen until the pressure becomes 1.1 MPa and temperature becomes 200℃. The work done during the process is 50 kJ. Find the heat transferred from the nitrogen to the surroundings. Use 𝑐𝑣 for nitrogen = 0.75 kJ/kg-K.


1
Expert's answer
2021-04-30T11:24:19-0400

"\nu=\frac mM=\frac{0.8}{0.028}=28.6~mol,"

"\gamma=1+\frac{R}{C_v}=1+\frac{8.31}{750}=1.011,"

"A=\frac{\nu R T_1}{\gamma -1}(1-(\frac{p_2}{p_1}\cdot \frac{T_1}{T_2})^{\gamma -1})=150~kJ,"

"Q=A-A'=100~kJ."